OpenStudy (kkutie7):

I got 178.15 my professor got 188.15 what is going on here?

OpenStudy (kkutie7):

\[\int\limits_{1}^{6}-.52x+37.45dx\]

OpenStudy (campbell_st):

well you get \[[-0.26x^2 + 37.45x]^6_{1}\]

OpenStudy (anonymous):

or \[-.26\times 6^2+.26+5\times 37.45\]

OpenStudy (campbell_st):

I'd say you're professor made a typo

OpenStudy (kkutie7):

Thank you =)